OpenStudy (anonymous):

2/3 of what number is 16

OpenStudy (wolfe8):

Well write the equation. Let's say the number is x. We have 2/3x = 16 Solve for x?

OpenStudy (anonymous):

24

OpenStudy (wolfe8):

... @madrockz I was going to guide him through it...

OpenStudy (john_es):

You still can do it.

OpenStudy (jack1):

ratios, so: \[\frac{ 2 }{ 3 } = \frac{ 16 }{ ??? }\] solve for ???

OpenStudy (jack1):

\[2\rightarrow 16 = \times 8 , so 3\rightarrow ???? = 3 \times 8\]

OpenStudy (anonymous):

@wolfe8 : Good on1...Bettr to guide instead of providing direct answer

OpenStudy (08surya):